# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

### v4

@@ 4.1.2. Opposition To AIGC
r/CryptoCurrency moderator: We’ve seen some ‘shills’ (aka companies or group of people) coming with AI content to promote their product…And of course promoting is detrimental, as it is often disguised as a ‘research’ post about that specific coin (for example), making it seem like a legit user researching a subject and that may influence buyers in comparison to a regular ad.
As we’ve just seen, our participants had a lot of opposition to AIGC use in their communities, but as we see next, some saw legitimate intentions behind its use.
@@ 4.1.3. Potential Benefits of AIGC
While the overall attitude that we heard from our participants towards AIGC in their communities was one of opposition, several participants did mention certain ways that AIGC could benefit their communities by increasing engagement. Participants’ attitudes towards the potential benefits of AIGC often came from what they perceived to be a legitimate, or at least good-faith, motivation behind AIGC posters. Five of our participants believed that a large amount of AIGC posting is motivated by a good-natured desire to participate. In particular, we heard from moderators who reporting thinking that community members were using AIGC to help complement their knowledge and writing ability:
@@
This participant was sympathetic with this motivation and engaged with the possibility that AIGC could be beneficial in certain situations.
In total we heard three separate moderators mention the potential for non-native English-speaking posters to use AI to improve their prose. These three moderators saw this as a legitimate desire to participate, but they emphasized that they wanted to hear the poster’s own voice, even if it was not perfect. We heard such a sentiment from a moderator of r/changemyview who shared: “I think AI as a tool is going to be useful…it’s useful for people who are poor writers, useful for people who aren’t strong English speakers. And when they need help, AI can fill in those gaps,” though they noted that it was important that such a contribution was still “sufficiently their words”. Of course, the definition of “sufficient” will be context dependent and may likely change over time as norms around AI use become more established.
One last attitude that we heard was that AIGC could be used to increase engagement in an otherwise inactive conversation. For example, one moderator speculated about the potential benefits of AIGC for platforms and provided a hypothetical example for us:
@@
Although our participants were aware of communities elsewhere on Reddit who believed that the potential benefits of AIGC outweighed the negatives, none of our participants shared this perspective. Partially due to our sampling strategy, our participants reported that their general attitude of opposition to AIGC in their communities overshadowed any potential benefits.
4.2. Findings: The Specific Perceived Threats of AIGC (RQ1)
Our participants’ general attitudes towards AIGC were often rooted in more specific threats that our participants believe AIGC poses to their communities. These perceived threats included lowering the quality of posts, straining relationships between community members, and amplifying coordinated attacks. We provide more detail about each of these threats below.
First, it is worth noting that our participants had different perceptions of the severity of the threats from AIGC, with some seeing it as severe while others regarded it as something that could easily be managed through moderation. When asked if AIGC was one of their top concerns, a moderator for r/CryptoCurrency noted that, “Early 2023, I would’ve said yes, but right now it has calmed down quite a bit.” On the other hand, some participants did regard AIGC as one of their highest concerns:
We see these two moderators as emblematic of a trend where our most concerned participants tended to moderate communities focused on authentic human connection, like r/explainlikeimfive, and the least concerned tended to moderate large subreddits that were less personal, such as r/CryptoCurrency. We will explore some of the detection challenges that the r/explainlikeimfive moderator references later in 4.4, but first we dig deeper into the why behind our participants’ different perceptions of the threat of AIGC to their communities.
4.2.1. Lowering Quality
Eleven moderators felt that posters using AIGC are “missing the point” of their communities. Subreddits exist for a wide range of reasons and adopt different values to meet their specific communities’ needs. We found that our participants assessed the quality of a contribution by considering how much it aligned with the values of their community. Our participants shared several specific ways that they thought AIGC lowered the quality of content in their communities, such as by decreasing the opportunities for authentic human connection, decreasing the reliability of information, and making conversations in the community more off-topic. All of these participants felt that low-quality content decreased the value that the community provides to its members and worried that it could result in decreased participation.
Less Human Connection
Seven participants mentioned the inherently inhuman nature of AIGC and expressed concern that this would lead to less authentic human connection in their communities. A moderator from r/changemyview summarized this sentiment well when they shared, “How can we change your view when you haven’t actually stated your view at all? This is ChatGPT’s view. And if I wanted to, I could go argue with ChatGPT. But that’s not helping here.” These participants believed that people come to their communities because they value connecting with other humans through dialogue and that AIGC does not meaningfully contribute to such conversations.
Less Accurate Information
Four participants who moderate communities that emphasize the accuracy of information (r/AskHistorians and r/explainlikeimfive) saw AIGC as threatening to their communities because AIGC is not necessarily factually sound. A moderator from r/AskHistorians shared that, “[inaccurate AIGC] undermines the whole intent of the sub, which is meant to give people that don’t have access…expertise in readable formats, from experts. And so, in the long term, it can undermine the trust that people have [in the community].” This participant described AIGC as “truthy-seeming” because of its tendency to include fake citations and felt that the appearance of being thoroughly researched increased the chances that readers would believe it.
More Off-Topic
Eleven participants said that they regarded AIGC as off-topic and felt that it would decrease the value of the community to its members. We heard several different reasons why our participants thought AIGC was off-topic. For example, six participants from communities that encourage or celebrate specific practices (e.g r/itookapicture, which is devoted to photography) said that they felt AIGC is off-topic because it is created without using their practice. When assessing the quality of a contribution to their community, these moderators were just as interested in the process used to create the contribution, as the contribution itself. The essence of their concern is captured by a moderator from r/WritingPrompts who noted, “[AIGC] is antithetical to everything the subreddit is about, which is about giving writers a chance to practice their craft and practice their skill.”
We also heard from five participants who thought that only original content was on-topic in their communities and that AIGC was not original content. For example, a participant who moderated r/WoW (a World of Warcraft subreddit where members frequently post fan art), said that, “On the art/other content side, we (on r/wow) consider it a matter of stealing w/o consent.” A moderator from r/AskHistorians echoed a similar sentiment when they described posting AIGC as, “taking words from something else. Even if it’s not a human, they are not your own words, and you are representing them as though they’re your own. So that’s plagiarism, right?”
4.2.2. Straining Relationships
In addition to feeling that AIGC was low-quality content, ten participants were concerned that AIGC would harm relationships in their communities. A moderator of a political subreddit was concerned about the general sentiment that emerged when legitimate users were falsely accused of posting AIGC. They noted that, “the possibility [of AIGC] fuels Redditors accusing other comments of being written by AI (which is a form of incivility).” A moderator of r/Zelda shared that the controversial issue of AIGC led to frequent off-topic arguments in their community. They reported that their main concern as a moderator was, “Do members feel safe and respected in our community,” and thus it was their their responsibility to remove, “arguments that shift away from civil debate towards petty fights / insults / trolling.” The participant saw that posts with AIGC often led to such arguments:
For this participant, banning AIGC from their community was thus a way to keep conversation civil.
4.2.3. Amplifying Attacks
Several participants mentioned the threat of AIGC being used to amplify malicious behavior like harassment and deception campaigns. However, moderators generally felt that existing techniques for countering these problems would be sufficient. For example, one participant shared concerns about AIGC being used as a tool for trolling, but they noted that they expected existing community defenses to be sufficient:
r/changemyview moderator: The one thing that I am worried about AI is that it does make trolling easier. You can generate larger amounts of trolling text and make it seem reasonable or seem like somebody wrote it for a lot longer…But we don’t tolerate trolling and actually it’s not that hard to tell when somebody’s trolling on our subreddit anyway. So it makes trolling easier [to do], but it’s something that I don’t think we’re going to be too stressed handling.
Another example is a moderator who shared about their experience with a ChatGPT-powered bot attack that was producing spam, which they were able to resolve with the help of Reddit admins:
r/AskHistorians moderator: A few months ago, we had an instance where we were subject to a bot attack that was using ChatGPT…Then we ended up reporting it to the admins who were able to take care of it through whatever they do on their back end.
In this example, the platform admins’ existing bot defenses were able to solve the problem, which made the participant view this kind of attack as a low concern.
The one exception to this was a moderator from r/news, which is one of the largest subreddits and frequently contains political content. This participant had seen their community targeted by, “a highly sophisticated deployment, an attempt to sway a narrative by utilizing accounts that had had their karma bolstered by automated commenting,” and were highly concerned that AIGC would make influence operations more frequent by facilitating karma-farming, which we discussed in 4.1.2 as an example of a bad-faith use of AIGC that our participants opposed. Our participant noted that while generative AI was much better at generating karma-worthy comments then earlier technologies, karma-farming was a threat only to the largest subreddits, which provide the greatest opportunity to gain karma.
While our participants reported different levels of concern about the threat that AIGC posed to their communities, all of them shared some degree of concern. Next, we share the changes that communities made to respond to these concerns.
@@ 4.3. Findings: Online Communities’ Responses to AIGC (RQ2)
Our participants shared various ways in which their communities have responded to concerns about AIGC, such as clarifying community norms and enforcing community rules.
@@ 4.3.1. Clarifying Community Norms
We heard from most participants that establishing a community stance towards AIGC was often a collaborative and iterative process between community members and moderators. Often it was the moderators who started the discussion, but other times it was the community:
@@
We heard from most participants that they tried to develop rules for AIGC in consultation with their wider community, not only moderators, to ensure that they were aligned with the community’s interests. One participant described how the community started off permissive towards AIGC before eventually voting to ban it altogether:
@@
This participant went on to describe the community’s process for publicly discussing issues, surfacing proposals for solutions, and soliciting votes on proposals from all community members—a process that their community used here to enact a rule banning AIGC.
All participants mentioned eventually codifying community norms into public subreddit rules. Some communities used rules that explicitly banned AIGC or bots, while others used more general rules (such as rules against plagiarism) as effective AIGC bans. Participants shared mixed results about the efficacy of these rules for decreasing the use of AIGC. We heard from seven participants that the volume of AIGC had decreased in response to enacting new rules, including from a moderator of r/AskHistorians who said, “People are using [AIGC] a little bit less now that they know what the rules are. They know that people don’t necessarily want it.” We also heard a moderator of r/explainlikeimfive express some skepticism about the effectiveness of rules: “I think it may stop some of the average users (to the extent that they actually read the rules, which is a reddit-wide problem.) Bots obviously don’t read rules.” This quote highlights that explicit rules can guide the behavior of good-faith users who want to act in accordance with the community’s values, but may not deter bad actors, especially if the rules are difficult to enforce.
@@ 4.3.2. Enforcing Community Rules
All of our participants’ communities devised—and tried to enforce—rules about AIGC. However, enforcement of these rules was often a challenge. We heard from three participants that needing to enforce rules against AIGC had increased their moderation workload:
@@
This imbalance between the amount of effort it takes to produce AIGC and the amount of effort that it takes to read it, especially since it tends to be stylistically more verbose than the average Reddit post, puts moderators at a disadvantage. We also heard from moderators who were concerned about alienating their community through false accusations. They noted that it took extra time and effort to get their decisions right, like a moderator of r/explainlikeimfive who shared, “You don’t want to be banning someone for using GPT when they don’t actually use it. So we might have to watch someone for a bit of time.” This need to spend time watching a community member for AIGC use is an unwelcome additional demand on the time of already over-burdened community moderators.
However, moderators were not the only ones involved in the rule enforcement process. Once a policy was in place, our participants relied on a combination of flagging by users, moderator review, automated tools, and platform action to enforce their rules. We heard from seven moderators that their community members often played an active role in rule enforcement, like a moderator of r/news who noted, “The first line of defense is the users themselves.” We also heard about instances where the load on the moderators became overwhelming, such as during the ChatGPT-powered bot-network attack discussed in 4.2.3, where moderators enlisted the help of Reddit platform admins to help them enforce their rules against bots.
Four participants mentioned that in addition to manual review, they used technical tools provided by the platform to enforce their rules, such as automoderator [^7] https://www.reddit.com/wiki/automoderator/, a programmable bot that scans and performs actions on all new posts. We also heard about additional tools available to the Reddit platform admins which were available for help in some severe cases of coordinated attacks:
@@
The primary challenge when enforcing rules about AIGC is the difficulty of detecting it. Next we will dig more deeply into the techniques that moderators reported using to aid them with detection.
4.4. Findings: Detecting AIGC (RQ2)
Our participants reported using both custom heuristics and technical solutions to detect AIGC in their communities. They tended to report that their detection techniques largely work now but project that they will cease to be effective as generative AI increases in sophistication:
@@ 4.4. Findings: Detecting AIGC (RQ2)
Below we discuss the challenges of detecting AIGC and provide examples of the heuristics and technical tools that moderators use to aid their work.
@@ 4.4.1. Detection Heuristics
Participants shared heuristics based on a variety of signals that they believe help them identify AIGC. To avoid aiding posters who are trying to evade these heuristics, we do not list the specifics of what our participants look for, but they broadly fell into the following categories: recognizable language patterns in posted content, details of users’ accounts and behavior, deviation from a community’s style norms, and inaccurate information.
@@
Twelve participants mentioned that certain signals in a post’s content, such as keywords, phrases, or distinct forms, may make moderators suspect that a post is AIGC. A moderator of r/Eldenring noted that AI-generated text is, “very repetitive i.e it tries to justify a point but instead of doing that, it ends up repeating the same point over and over to the point of ad nauseum.” A moderator of r/itookapicture shared that with AI-generated images, “there is often something that feels a bit unnatural…The images look like the average of something, rather than a unique individual with all of the flaws that come with that.”
@@
Six participants reported that they often consider the details of a poster’s account, including their past posting behavior, when they suspect that a post is AIGC. One red flag that alerted participants was a dramatic change in the style of a poster’s prose. For example, a moderator of r/changemyview shared that suspicious posts were, “often very different in terms of writing from what the user posted in response to other users questions or comments.”
We heard from four participants that signals used to detect more traditional bots may also be relevant when identifying AIGC, like a moderator of r/Zelda who noted that, “accounts used for spam…there are other indicators to find those (account age, history,…) Even before ChatGPT, we would find copy-pasted or markov-generated comments with these same account indicators.”
@@
Four participants reported using their knowledge of their community’s particular communication norms to identify posts that don’t fit as AIGC:
@@ 4.4.2. Technical Tools
Though they acknowledged their unreliability, most of our participants reported using a variety of technical tools in detection efforts, such as black-box detection tools provided by third parties, and community-developed solutions. Participants mentioned programming the detection heuristics referenced above into automated tools like automoderator, which then apply the heuristics to every new post and flag anything suspect. Six participants mentioned manually copying posts into third-party detection tools to see if they were AIGC. A moderator of r/writingprompts reported that they had tested, but never deployed, a homespun bot that would apply third-party detectors automatically to all new posts.
At the same time, a moderator of r/changemyview mentioned that because they are aware of the unreliability of third-party detection tools, they rarely defer to them absolutely: “I don’t necessarily trust the tool–that’s why we have several layers.…We always try to talk to the user…But still, I think the tool is useful.” Two participants mentioned that the tools were least effective on shorter pieces of text, like a moderator of r/writingprompts who noted that the detector they were using, “was very bad with short comments. If a comment was like 100 words or a really short story–every poem got flagged as AI. And like, No, it’s not. It’s just short.”
@@ 5. Discussion
Our qualitative investigation offers two main research contributions: (1) a description of volunteer community moderators attitudes towards AIGC use in their communities and (2) a description of the moderators’ and their online communities’ responses to AIGC and the challenges that they’ve encountered governing its use. We place our work in dialog with Seering’s call to the CSCW research community to center community self-moderation (Seering20) and specifically see our study as motivated by the second “guiding question” proposed by Seering: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” We believe that allowing communities to make context-sensitive decisions about AIGC is central to their self-determination and should be an important issue for the CSCW research community. Below we discuss the implications of our findings and potential ways that our research may be able to aid online communities during this period of technological change.

### v5

@@ 4.1.3. Governance Concerns
r/CryptoCurrency moderator: We’ve seen some ‘shills’ (aka companies or group of people) coming with AI content to promote their product…And of course promoting is detrimental, as it is often disguised as a ‘research’ post about that specific coin (for example), making it seem like a legit user researching a subject and that might influence buyers in comparison to a regular ad.
In addition to broad attacks on a community, we also heard from participants who were concerned about AIGC being used for targeted attacks on individual members. One moderator of r/politics reported that they associated AIGC with, “usually some form of trolling, or ‘outsourcing’ writing a comment the user wanted to make.” Another participant also mentioned trolling, though they said that existing moderation solutions for this problem would be sufficient:
r/changemyview moderator: The one thing that I am worried about AI is that it does make trolling easier. You can generate larger amounts of trolling text and make it seem reasonable or seem like somebody wrote it for a lot longer…But…it’s not that hard to tell when somebody’s trolling on our subreddit anyway. So it makes trolling easier [to do], but it’s something that I don’t think we’re going to be too stressed handling.
Finally, we heard concerns about AI making it easier to post objectionable content. One participant who moderated r/Zelda shared that they had seen AI used to produce specific types of objectionable content that their community had already banned, such as commercial or NSFW content. They noted that while this type of content was not new, “it seemed to be new territory for people to expedite these things via AI.”
@@ 4.1.4. Potential Benefits of AIGC
While our recruiting approach and research questions primarily surfaced an overall attitude of opposition, several moderators mentioned ways that AIGC could potentially benefit their communities by increasing engagement. These moderators perceived a legitimate, or at least good-faith, motivation behind posters who used AIGC to complement their knowledge and writing ability:
@@
Three moderators mentioned that AI could help non-native English-speakers improve their posts. We heard from a moderator of r/changemyview who thought that, “AI as a tool is going to be useful…it’s useful for people who are poor writers, useful for people who aren’t strong English speakers. And when they need help, AI can fill in those gaps,” though they noted that it was important that such a contribution was still, “sufficiently their words”. Of course, the definition of “sufficient” will be context dependent and may change over time as norms around AI use become more established.
One last benefit that we heard was that AIGC could increase engagement in an inactive conversation. For example, one moderator speculated about the potential benefits of AIGC for platforms and provided a hypothetical example:
@@
On the whole, our evaluation of the interview transcripts suggests that despite these potential benefits, all of our participants felt that, at least for now, the negative aspects of AIGC outweigh the good.
@@ 4.2. Findings: Online Communities’ Responses to AIGC (RQ2)
To counter the perceived threats of AIGC, moderators have enacted specific changes in their communities. First, our participants emphasized collaborating with their community members to clarify norms about the use of AIGC. Once clarified, these norms could be codified into rules, which moderators were then tasked with enforcing.
@@ 4.2.1. Clarifying Community Norms
Most participants spoke of establishing a community stance towards AIGC through a collaborative and iterative process involving community members and moderators. Often it was the moderators who started the discussion, but other times it was the community:
@@
Most participants reported they wanted to ensure that their stance on AIGC was aligned with their communities’ interests. One participant said that their community initially allowed AIGC before eventually voting to ban it:
@@
This participant went on to describe their community’s process for discussing issues, surfacing proposals for solutions, and soliciting proposal votes from all community members—a process that they used here to enact a rule banning AIGC.
Almost all participants mentioned eventually codifying community norms into public subreddit rules. Some communities used rules that explicitly banned AIGC or bots, while others used more general rules (such as rules against plagiarism) as effective AIGC bans. Participants shared mixed results about the efficacy of these rules. We heard from seven participants that the volume of AIGC had decreased in response to new rules, including from a moderator of r/AskHistorians who said, “People are using [AIGC] a little bit less now that they know what the rules are. They know that people don’t necessarily want it.” However, a moderator of r/explainlikeimfive expressed skepticism about the efficacy of rules: “I think it may stop some of the average users (to the extent that they actually read the rules, which is a reddit-wide problem.) Bots obviously don’t read rules.” As this participant suggested, explicit rules can guide the behavior of good-faith users, but might not deter bad actors.
@@ 4.2.2. Enforcing Community Rules
All of our participants’ communities devised—and tried to enforce—rules about AIGC. However, the difficulty of detecting AIGC meant that enforcement of such rules was often challenging. Three moderators spoke of AIGC increasing their workload:
@@
This imbalance between the effort to produce AIGC and to read it, given AIGC tends to be verbose, puts moderators at a disadvantage. Also, given how alienating a false accusation can be, moderators must take extra care with their decisions; a moderator of r/explainlikeimfive shared, “You don’t want to be banning someone for using GPT when they don’t actually use it. So we might have to watch someone for a bit of time.”
Our participants used a variety of strategies to try to make enforcement of these rules less burdensome. Seven moderators mentioned sharing the work with community members, like a moderator of r/news who noted, “The first line of defense is the users themselves.” However, this runs the risk of causing incivility between members in the case of false accusations, as mentioned in Section 4.1.2.
Four moderators spoke of their reliance on platform tools, especially automoderator [^7] https://www.reddit.com/wiki/automoderator/, a programmable bot that scans and performs actions on all new posts. Such automated tools lessened the manual effort required by moderators, but they are only as effective as the detection heuristics programmed into them, which, as we discuss in the next section, can be flawed. We heard from moderators that such tools are not enough to catch all AIGC, and in cases of extreme post volume, such as during coordinated bot attacks, moderators sometimes sought help from Reddit platform admins:
@@
All of these enforcement techniques ultimately require making decisions about what is and is not AIGC. Next we discuss our participants’ reported methods for making these difficult decisions.
4.3. Findings: Methods for Detecting AIGC (RQ2)
Our participants shared both custom heuristics and technical solutions that they use to detect AIGC. Even if participants thought their detection techniques mostly work now, they worry that their tools will be less effective as generative AI improves:
@@ 4.3.1. Detection Heuristics
There are a number of signals that moderators believe help them identify AIGC. To avoid aiding bad-faith actors, we do not list specific tells, but they broadly fell into the following categories: recognizable language patterns in posted content, details of users’ accounts and behavior, deviation from a community’s style norms, and inaccurate information.
@@
Twelve participants mentioned that certain patterns in a post’s content, such as keywords, phrases, or distinct forms, may make moderators suspect that a post is AIGC. A moderator of r/Eldenring noted that AI-generated text is, “very repetitive i.e it tries to justify a point but instead of doing that, it ends up repeating the same point over and over to the point of ad nauseum.” A moderator of r/itookapicture shared that with AI-generated images, “there is often something that feels a bit unnatural…The images look like the average of something, rather than a unique individual with all of the flaws that come with that.”
@@
When confronted with suspect content, six participants reported that they often consider the details of a poster’s account, including their past posting behavior. One red flag is a dramatic change in the style of a poster’s prose. For example, a moderator of r/changemyview shared that suspicious posts were, “often very different in terms of writing from what the user posted in response to other users questions or comments.” Additionally, four moderators, like a moderator of r/Zelda, noted that signals that identify more traditional bots are also relevant for identifying AIGC: “accounts used for spam…there are other indicators to find those (account age, history,…) Even before ChatGPT, we would find copy-pasted or markov-generated comments with these same account indicators.”
@@
Four participants reported using their knowledge of their communities’ particular communication norms to identify atypical posts as AIGC:
@@ 4.3.2. Technical Tools
Though they acknowledged such tools’ unreliability, most participants used automated approaches to detect AIGC, including both third-party, black-box detection tools and community-developed solutions. Participants spoke of programming the detection heuristics mentioned above into tools such as automoderator, which apply the heuristics to every new post and flag anything suspect. Six participants mentioned manually copying posts into third-party detection tools to see if they were AIGC. A moderator of r/writingprompts reported that they had tested, but never deployed, a homespun bot that would apply third-party detectors automatically to all new posts.
At the same time, a moderator of r/changemyview mentioned that because they are aware of the unreliability of third-party detection tools, they rarely defer to them absolutely: “I don’t necessarily trust the tool–that’s why we have several layers.…We always try to talk to the user.” Two participants mentioned that the tools were least effective on shorter text, like a moderator of r/writingprompts who noted that one tool, “was very bad with short comments. If a comment was like 100 words or a really short story–every poem got flagged as AI. And like, No, it’s not. It’s just short.”
@@ 5. Discussion
Our qualitative investigation offers two main research contributions: (1) a description of moderators’ attitudes towards AIGC in subreddits that restrict its use and (2) insights into how these moderators and their communities have responded to the challenges of AIGC. On the whole, our findings align with those of recent studies into the impact of generative AI on online communities (burtch_consequences_2024; wei2024understandingimpactaigenerated), which emphasize that the social value of online communities is more important than ever in the age of generative AI. Platform designers and community stewards should focus on ways to increase this social value, and one way to do this is by empowering communities to take their own stance on AIGC.
As generative AI usage grows, we place our work in dialog with Seering’s call to the CSCW research community to center community self-moderation (Seering20), especially their second “guiding question”: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” We believe our findings demonstrate that allowing communities to make context-sensitive decisions about AIGC is central to their self-determination and should be an important topic for CSCW research. To offer implications for design, we expand below on the values that shape community attitudes towards AIGC as well the importance of clear and enforceable community norms.